Assume that the average annual salary for a worker in the United States is $30,000 and that the annual salaries for American are normally distributed with a standard deviation equal to $7,000. Find the following:
a)What percentage of Americans earn below $22,000?
b)What percentage of Americans earn above $42,000?

Answer by stanbon(75887)

Assume that the average annual salary for a worker in the United States is $30,000 and that the annual salaries for American are normally distributed with a standard deviation equal to $7,000. Find the following:
a)What percentage of Americans earn below $22,000?
z(22000) = (22000-30000)/7000 = -8/7
-------------------------------------------
P(x < 22000) = P(z < -8/7) = 0.1265
-----------------------------------------------
b)What percentage of Americans earn above $42,000?
z(42000) = (42000-30000)/7000 = 1.7143
P(x > 42000) = P(z > 1.7143) = 0.0432
===========================================
Cheers,
Stan H.